# Key Words.: binaries: spectroscopic – binaries: close – stars: early-type – stars: massive – Magellanic Clouds1 Introduction

### v1

@@ 7 Conclusions
High Multiplicity fraction at Low Metallicity: We detect 153 spectroscopic binaries (91 SB1, 59 SB2, 3 SB3) based on significant RV variability. This corresponds to an observed multiplicity fraction of % , exceeding reported values for B-type stars in the MW and LMC, and suggesting a potentially larger multiplicity fraction at lower metallicities.
@@
Intrinsic multiplicity fraction from independent methods: After correcting for observational biases with MC simulations, we infer an intrinsic close-binary fraction of % . For this, we consider as binaries those systems which simultaneously fulfil having peak-to-peak RV variations above a threshold of km s-1 and an RV significance above , consistent with previous studies of massive stars in the MW and LMC. Experimenting with a higher RV threshold reduces the multiplicity fraction down to % for km s-1. However, an independent MCMC analysis which does not rely on binary criteria but considers the full distribution returns a multiplicity fraction of % . Letting the distribution of orbital periods to vary, further confirms the high multiplicity fraction, resulting in % , although the orbital period distribution cannot be well constrained with our data.
@@
Potential anti-correlation with metallicity: The resulting multiplicity fraction is significantly higher in comparison to similar samples in the LMC and MW. This is consistent with an anti-correlation between metallicity and close-binary fraction among B-type III/V stars. While the fraction of O-type binaries seems to remain constant across different metallicities, B-type stars might be affected by how soon disk fragmentation occurs during massive star formation. B-type binaries forming during extremely early fragmentation would have enough time and disk material to spiral inwards to very short orbits ( d). At later times however, low-metallicity disks have had more time to cool down due to the reduced opacity, increasing fragmentation and therefore multiplicity. However, with less accreting material, the star is not able to migrate to such low separations. This would imply that the increased multiplicity fraction at low metallicity is dominated by binaries with orbital periods above 10 d, while the fraction of shorter-period binaries should remain invariant with metallicity, which is consistent with previous studies of eclipsing binaries.
@@
Implications for the early Universe: The SMC’s metallicity is similar to that of many high-redshift, star-forming galaxies. A higher fraction of close binaries at low metallicity would translate to more frequent interactions, potentially increasing the number of exotic transients, and envelope-stripped stars, relevant for cosmic reionisation. Our results also provide important input for population synthesis studies at low metallicity, and are critical for the population of gravitational-wave progenitors.

### v4

@@ 7 Conclusions
High multiplicity fraction at low metallicity: We detect 153 spectroscopic binaries (91 SB1, 59 SB2, 3 SB3) based on significant RV variability. For this, we consider as binaries those systems which simultaneously fulfil having peak-to-peak RV variations above a threshold of km s-1 and an RV significance above , consistent with previous studies of massive stars in the MW and LMC. This corresponds to an observed multiplicity fraction of % , exceeding reported values for B-type stars in the MW and LMC, and suggesting a potentially larger multiplicity fraction at lower metallicities.
@@
Intrinsic multiplicity fraction from independent methods: After correcting for observational biases with MC simulations, we infer an intrinsic close-binary fraction of % . Experimenting with a higher RV threshold reduces the multiplicity fraction down to % for km s-1. However, an independent MCMC analysis, which does not rely on binary criteria but considers the full distribution, returns a binary fraction of % , confirming the high multiplicity. Letting the distribution of orbital periods vary further reaffirms our result, leading to % , although the orbital period distribution cannot be well constrained with our data.
@@
Potential anti-correlation with metallicity: The resulting multiplicity fraction is significantly higher in comparison to similar samples in the LMC and MW. This is consistent with an anti-correlation between metallicity and close-binary fraction among B-type III/V stars. While the fraction of O-type binaries seems to remain constant across different metallicities, B-type stars might be affected by how early or late disc fragmentation occurs during massive star formation. B-type binaries forming during extremely early fragmentation would have enough time and disc material to spiral inwards to very short orbits ( d). At later times however, low-metallicity discs have had more time to cool down due to the reduced opacity, increasing fragmentation and therefore enhancing multiplicity. However, with less accreting material, the star is not able to migrate to such low separations. This implies that the increased multiplicity fraction at low metallicity predominantly involves binaries with orbital periods above 10 d, while the fraction of shorter-period binaries remains invariant with metallicity, consistent with previous studies of EBs.
@@
Implications for the early Universe: The SMC’s metallicity is similar to that of many high-redshift, star-forming galaxies. A higher fraction of close binaries at low metallicity would translate to more frequent interactions, potentially increasing the number of exotic transients and envelope-stripped stars, relevant for cosmic reionisation. Our results also provide important input for population synthesis studies at low metallicity, and are critical for the population of gravitational-wave progenitors.
Data availability
Tables D.1 and D.2 are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ Acknowledgements.
- Hastings et al. (2021) Hastings, B., Langer, N., Wang, C., Schootemeijer, A., & Milone, A. P. 2021, A&A, 653, A144